OpenStudy (anonymous):

Please help me complete these identities: 1) cscx cotx / secx 2) sec^2x csc^2x

OpenStudy (anonymous):

rewrite all in terms of sine and cosine and it will be much easier

OpenStudy (anonymous):

I did and I'm stuck... 1) (1/sinx 1/tanx)/ 1/cosx 2) 1/cos^2x (1/sin^2x) Where do I go from here? I've been up most of the night working on this kind of problems and my brain is beyond fried.

OpenStudy (anonymous):

lets replace cosine by \(a\) and sine by \(b\) then the first one is \[\frac{\frac{1}{b}\times \frac{b}{a}}{\frac{1}{a}}\]

OpenStudy (anonymous):

nothing left but algebra you get \[\frac{b}{ab}\times \frac{a}{1}\] \[=\frac{ab}{ab}=1\]

OpenStudy (anonymous):

oh damn i made a mistake, scratch that

OpenStudy (anonymous):

you actually get \[\frac{\frac{1}{b}\times \frac{a}{b}}{\frac{1}{a}}\] \[=\frac{a}{b^2}\times \frac{a}{1}=\frac{a^2}{b^2}\]

OpenStudy (anonymous):

since\(a\) is cosine and \(b\) is sine, this is \(\frac{\cos^2(x)}{\sin^2(x)}=\cot^2(x)\)

OpenStudy (anonymous):

I worked the first one out and i got cot^2x. Oh my gosh I completely forgot about cos/sin. Thank you!!!

OpenStudy (anonymous):

How would I go about finishing the next one?